OpenStudy (lgbasallote):

Universal set \(\sigma\) = {7, 8, 9, 10, 11, 12, 13, 14, 15} A = {numbers that are multiples of 3} B = {odd numbers} C = {10 \(\le\) number \(\le\) 13} how do I know how many numbers there are in set \(A \cup B\)

OpenStudy (anonymous):

AUB = { 7,9,11,12,13,15}

OpenStudy (lgbasallote):

???

OpenStudy (anonymous):

A={9,12,15} B = {7,9,11,13,15}

hartnn (hartnn):

n(A)= 3 n(B)= 5 n(A intersection B) = 2 (9 and 15) so n(A union B) = n(A)+n(B)- n(A intersection B) = 3+5-2 = 6

hartnn (hartnn):

u got that lg ?

OpenStudy (lgbasallote):

less math more words?

OpenStudy (anonymous):

hartnn's way is a analatical thing.

OpenStudy (anonymous):

it is more easy to draw a diagram..

OpenStudy (lgbasallote):

aren't there infinite numbers in set A U B?

OpenStudy (anonymous):

i think this can be more difficult for you to solve as a way of hartuun

OpenStudy (lgbasallote):

like i said before...less math more words....i really am not a ver good math learner

OpenStudy (lgbasallote):

in case you haven't noticed my "i hate math" [picture

hartnn (hartnn):

the universal set here is finite so how come union ever become infinite ?

OpenStudy (lgbasallote):

oh

OpenStudy (lgbasallote):

the universal set limits the elements

OpenStudy (lgbasallote):

you could have said that earlier...

OpenStudy (anonymous):

you said.. universal set is cofined in several number//

OpenStudy (anonymous):

i thought you noticed..

hartnn (hartnn):

yup, we see w.r.t universal set mentioned/given.

OpenStudy (lgbasallote):

i also said i didn't know what those meant

OpenStudy (lgbasallote):

so "numbers multiple of three" are the ones in the universal set?

OpenStudy (anonymous):

universal set means that total group is that number.

OpenStudy (anonymous):